actice final exam Saved Montgomery Industries spent $660,000 in 2015 on a construction project to build a Montgomery also capitalized $33,000 of interest on the project in 2015. Montgomey man ced 100% ofthe construction with a 10% construction loan. The project was completed on September 30, 2016. Additional expenditures in 2016 were as follows: Feb. 28 Apr. 30 Jul. 1 Sept. 30 $95,400 185,400 41400 69,400 Required: Determine the completed cost of the library. (Do not round intermediate calculations.) cost of the library 1,091,530
Answer #1
$ Period Average expenditure
Accumulated expenditure Jan 1 693000 9/9 693000
Feb. 28 95400 7/9 74,200
Apr. 30 185400 5/9 103,000
Jul. 1 41400 3/9 13,800
Sept. 30 69400 0 0
Average Accumulated expenditure 1,084,600 884,000
Interest to be capitalized = 884000*10%*9/12= $66,300
Completed lost of Library = 1084600+66300= $1,150,900

IN CASE OF ANY DOUBTS OR CORRECTIONS FEEL FREE TO COMMENT BELOW

PLEASE RATE MY ANSWER BY HITTING ??

THANK YOU